Find 76.5 ÷ 0.15 Find 76.5 ÷ 0.15 Find 76.5 ÷ 0.15 Find 76.5 ÷ 0.15

Find 
76.5 ÷ 0.15
Find 
76.5 ÷ 0.15
Find 
76.5 ÷ 0.15
Find 
76.5 ÷ 0.15

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

We have,
= (765/10) ÷ (15/100)
= (765/10) × (100/15)
= (765 × 100)/ (10 × 15)
= (51 × 10)/ (1 × 1)
= 510

Thanks